# Three-Dimensional Almost Contact Metric Manifolds Revisited via the Newman-Penrose Formalism

### v1

@@ 2 Preliminary
$=\kappa\text{Ric}(\xi,\xi)+(\epsilon-2\bar{\rho}-\rho)\text{Ric}(\xi,\partial)$ (2.17)
$+\sigma\text{Ric}(\partial,\overline{\partial})-(\overline{\kappa}+2\overline{\beta})\text{Ric}(\partial,\partial)-\kappa\text{Ric}(\partial,\overline{\partial}),$ (2.18)
$\partial(\text{Ric}(\xi,\overline{\partial}))+\overline{\partial}(\text{Ric}(\xi,\partial))-\xi(\text{Ric}(\partial,\overline{\partial}))-\frac{1}{2}\xi(\text{Ric}(\xi,\xi))$ (2.19)
$=-(\rho+\overline{\rho})(\text{Ric}(\xi,\xi)-\text{Ric}(\partial,\partial))$ (2.20)
$-\overline{\sigma}\text{Ric}(\partial,\partial)-\sigma\text{Ric}(\partial,\overline{\partial})-(2\overline{\kappa}+\overline{\beta})\text{Ric}(\xi,\partial)-(2\kappa+\beta)\text{Ric}(\partial,\overline{\partial}).$ (2.21)
@@ 4 The $\eta$-Einstein condition on three-dimensional ACM manifolds
There exists a non-compact and non-complete rank $3$ normal ACM manifold which is $\eta$-Einstein but neither $\alpha$-Sasakian nor $\beta$-Kenmotsu. For instance, consider the manifold $M=\mathbb{R}^{3}\setminus\{(x,y,0)\mid y\in\pi\mathbb{Z}\}$ equipped with the metric
$g=\frac{1}{2}(z^{2}+\sin^{2}y)(dx^{2}+dy^{2})+(dz+\cos(y)dx)^{2},$ (4.42)
$\eta=dz+\cos(y)dx.$ (4.43)
Remark 4.6.
@@ 5 Conclusions and Discussion
Finally, we comment on the classification of three-dimensional normal ACM $\eta$-Einstein manifolds. Although we have provided a classification for the compact case in this study, the non-compact case has been left untouched. Since, in the case of rank 1, the construction of $\beta$-Kenmotsu $\eta$-Einstein structures in the non-compact setting is analogous to the compact one, the non-trivial case of interest is rank 3. By using the Raychaudhuri equations obtained from the generalized Sachs equation,
$\xi(\Theta)=\Theta^{2}-\omega^{2}+\frac{1}{2}\textrm{Ric}(\xi,\xi),\quad-\xi(\omega)=2\Theta\omega,$
the behavior of $\omega$ and $\Theta$ along the Reeb orbits can be fully described in terms of elementary functions. However, a major difficulty arises in the non-compact case regarding the control of these quantities in the horizontal directions. In particular, when $\inf\omega=0$ or the $\varphi$-sectional curvature is not bounded below, $\eta$-Einstein solutions are expected to exhibit complex behavior even in the complete case.

### v2

@@ 2 Preliminary
$=\kappa\text{Ric}(\xi,\xi)+(\epsilon-2\rho-\bar{\rho})\text{Ric}(\xi,\partial)$ (2.17)
$+\sigma\text{Ric}(\xi,\overline{\partial})-(\overline{\kappa}+2\overline{\beta})\text{Ric}(\partial,\partial)-\kappa\text{Ric}(\partial,\overline{\partial}),$ (2.18)
$\partial(\text{Ric}(\xi,\overline{\partial}))+\overline{\partial}(\text{Ric}(\xi,\partial))-\xi(\text{Ric}(\partial,\overline{\partial}))+\frac{1}{2}\xi(\text{Ric}(\xi,\xi))$ (2.19)
$=-(\rho+\overline{\rho})(\text{Ric}(\xi,\xi)-\text{Ric}(\partial,\bar{\partial}))$ (2.20)
$-\overline{\sigma}\text{Ric}(\partial,\partial)-\sigma\text{Ric}(\overline{\partial},\overline{\partial})-(2\overline{\kappa}+\overline{\beta})\text{Ric}(\xi,\partial)-(2\kappa+\beta)\text{Ric}(\xi,\overline{\partial}).$ (2.21)